OpenStudy (anonymous):

Find the equation for the parabola with focus at (3, -3) and vertex at (3, 1). a. x^2 + 6x +16y -7 = 0 b. x^2 + 6x + 16y - 25 = 0 c. x^2 - 6x - 16y + 25 = 0 d. x^2 - 6x + 16y -7 = 0

OpenStudy (anonymous):

Vertex (3,1) and Focus (3,-3) From these two we can deduce that the axis of the parabola is parallel to the x-axis.  Let, \[X ^{2}=4aY\] be the eq. of the parabola. \[X=(x-3) \] \[Y=(y-1)\] So, in real axis, the equation of the parabola is, \[(x-3)^{2}=4a(y-1)\] Now, a=Distance between the vertex and the focus=\[\sqrt{(3-3)^{2} + (1-(-3))^{2}} = 4\] Therefore the equation of the parabola is , \[(x-3)^{2}=16(y-1)\] Solving which will give you, \[x ^{2}-6x-16y+25=0\] Therefore, (c) is the answer.